# Knowledge Enhanced Conditional Imputation for Healthcare Time-series

### v3

@@ 5.3 Experimental Results
[…]
Weighting Parameter Comparison
The non-uniform masking probability for a given feature is determined the parameters $U$ and $I$, in addition to the feature’s prior probability as shown in Eq.(21)- (22). Having examined CSAI’s performance with different masking ratios (i.e $U$ values) in Section 5.3, we now study the effect of the weighting parameter $I$ on the resulting imputation and classification performance. The findings are shown in Table 4 and are further illustrated in Fig.4. An optimal weighting parameter, shown to be around 5, results in the lowest imputation error, suggesting that a balanced representation of features is crucial for accuracy. However, for classification, increasing the weighting parameter leads to higher errors and a marginal decrease in the AUC, highlighting that excessive weighting may not uniformly improve performance across different machine learning tasks. These findings reveal the interplay between feature representation adjustments and task-specific model efficacy, underscoring the importance of calibration in the non-uniform masking approach for complex time series data.
Table 4: Impact of weighting parameter $I$ on imputation and classification performance metrics. The table showcases the number of epochs required, imputation MAE and classification AUC across varying weighting factors.
@@
6 Discussion and Conclusions
This study introduced CSAI, a novel approach for imputing and classifying missing data in multivariate medical time series. CSAI demonstrated significant potential to improve the accuracy and reliability of medical data analysis, which is critical for high-stakes healthcare decision making. Using conditional knowledge embedding and attention mechanisms, CSAI outperformed established benchmarks in multiple real-world healthcare datasets, particularly in scenarios with lower masking ratios. Its ability to model non-random missingness and capture complex temporal relationships positions it as a valuable tool for real-world clinical applications.
Our comprehensive experimental evaluation showcased the efficacy of CSAI in diverse healthcare datasets, where it frequently surpassed baseline state-of-the-art models such as BRITS, BRITS-GRU, and others. CSAI consistently achieved the lowest imputation error and competitive classification performance in various masking configurations. The strong results suggest that CSAI can be instrumental in providing a strong base for predictive models used in various downstream tasks, including early diagnosis, deterioration prediction, and treatment recommendations in healthcare environments where incomplete data is a common challenge.
@@ 6 Discussion and Conclusions
Despite these promising results, there are areas where CSAI’s performance indicates room for improvement. In the MIMIC_89f dataset, CSAI slightly underperformed compared to the baseline, highlighting potential limitations in handling datasets with a higher proportion of categorical variables. This suggests a need for further refinement in the integration of domain-specific knowledge into the model and additional exploration of training data strategies to improve its robustness.
@@ 7 Acknowledgments
8 Data availability
The datasets generated and/or analysed during the current study are available from the corresponding authors on reasonable request.
9 Ethics declarations
9.1 Conflict of Interests
The authors declare that they have no conflict of interest with respect to the publication of this paper.

### v4

@@ V-D Experimental Results
The non-uniform masking probability used by CSAI for each feature is determined the parameters $U$ and $I$, in addition to the feature’s prior probability as shown in Eq.(21)- (22). We studied the effect of the weighting parameter $I$ on the resulting imputation and classification performance. The findings are shown in Table VI and are further illustrated in Fig.5. An optimal weighting parameter, shown to be around 5, results in the lowest imputation error, suggesting that a balanced representation of features is crucial for accuracy. However, for classification, increasing the weighting parameter leads to higher errors and a marginal decrease in the AUC, highlighting that excessive weighting may not uniformly improve performance across different machine learning tasks. These findings reveal the interplay between feature representation adjustments and task-specific model efficacy, underscoring the importance of calibration in the non-uniform masking approach for complex time series data.
TABLE VI: Impact of weighting parameter $I$ on imputation and classification performance metrics. The table showcases the number of epochs required, imputation MAE and classification AUC across varying weighting factors.
@@
VI Discussion and Conclusions
This study introduced CSAI, a novel approach for imputing and classifying missing data in multivariate medical time series. CSAI’s novelty lies in the provision of components specifically tailored to medical time-series, where the frequency and timing of data collection is highly variant, and long- and short-term correlations are pervasive. CSAI demonstrated significant potential to improve the accuracy and reliability of medical data analysis, which is critical for high-stakes healthcare decision making. Using conditional knowledge embedding and attention mechanisms, CSAI outperformed established benchmarks in multiple real-world healthcare datasets, particularly in scenarios with lower masking ratios. Its ability to model non-random missingness and capture complex temporal relationships positions it as a valuable tool for the clinical domain.
Our comprehensive experimental evaluation showcased the efficacy of CSAI in diverse healthcare datasets, where it frequently surpassed baseline state-of-the-art models such as BRITS, BRITS-GRU, and others. CSAI consistently achieved the lowest imputation error and competitive classification performance in various masking configurations. The strong results suggest that CSAI can be instrumental in providing a strong base for predictive models used in various downstream tasks, including diagnosis and outcome prediction.
@@ VII Acknowledgments
VIII Availability
CSAI is an open-source model. All code and experimental details are available on the CSAI GitHub repository https://github.com/LinglongQian/CSAI. CSAI is integrated into the PyPOTS ecosystem https://docs.pypots.com/en/latest/pypots.imputation.html#module-pypots.imputation.csai and is immediately callable from the Python package.